OpenStudy (anonymous):

A boat travels 24 miles upstream in 3 hours, but when going downstream it is able to cover 36 miles in 3 hours. What is the speed of the boat in still water and the speed of the current?

OpenStudy (anonymous):

Let b and c be the speed of the boat and current respectively. Rate equals distance divided by time, r = d/t. Solve the following simultaneous equations for b and c.\[\left\{b-c=\frac{24}{3},b+c=\frac{36}{3}\right\} \]\[\{b=10 \text{ mph},c=2 \text{ mph}\} \]